Session: Jun 2013 (session 2)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (C.Bortolotti at Fri May 24 10:09:13 2013) Success:
No known problems

Torun (K. Borkowski at Fri May 24 10:15:30 2013) Success:
No known problems.

Yebes (Daniel Valero at Fri May 24 10:21:53 2013) Success:
No known problems

Hartebeesthoek (Jonathan Quick at Fri May 24 10:46:57 2013) Success:
No known problems

Svetloe (Isaenko. at Fri May 24 10:52:30 2013) Minor failures: RECO
Scans No0007-No0011,No0013-No0017 failed due to Mark5 disk pack problem.

Zelenchukskaya (Lysenkova M. at Fri May 24 10:52:50 2013) Success:

Badary (Sorokovikov at Fri May 24 10:54:55 2013) Minor failures:
Scan No0008 skipped due to power cut

Urumqi (Peng Li at Fri May 24 11:56:56 2013) Success:
Observing only the last 40 minutes because of other tasks

Noto (L. Nicotra at Fri May 24 12:13:47 2013) Success:
No problem.

Onsala (Michael Lindqvist at Sun May 26 13:19:47 2013) Minor failures: TSYS
Used both the DBBC and the old rack. Problems with Tsys in DBBC data, it is ok for the old rack. The reason is that we forgot to turn off Tsys-measurments for the old rack.

Effelsberg (Georgi at Mon May 27 08:02:36 2013) Success:

Westerbork (Geert Kuper at Mon May 27 19:11:26 2013) Success:

Shanghai (BoXia at Fri May 31 06:05:34 2013) Success:

Irbene (Vladislavs Bezrukovs at Fri Jul 05 09:06:21 2013) Success:
not known problems